A function f(x) satisfies f(x+y)=f(x)+y for all x, y belongs to R and f(0)=5. Then f(2020) is equal to

Grade:12

1 Answers

Soumendu Majumdar
5 years ago
Dear Student,
f(2020) can be written as
f ( 0 + 2020 )
= f ( 0 )  + 2020    since f ( x + y ) = f ( x ) + y
= 5 + 2020     since f ( 0 ) = 5
= 2025
Hope it helps!